A particular manufacturing design requires a shaft with a diameter of 18.000 mm, but shafts with diameters between 17.987 mm and 18.013 mm are acceptable. The manufacturing process yields shafts with diameters normally​ distributed, with a mean of 18.005 mm and a standard deviation of 0.004 mm. Complete parts​ (a) through​(d) below.

a. For this​ process, what is the proportion of shafts with a diameter between 17.987 mm and 18.000 mm​?

b. For this process, what is the probability that the shaft is acceptable?

c. For this process what is the diameter that will be exceeded by only 5% of the shafts?

d. What would your answers to parts (a) through (c) if the standard deviation of the shaft diameters were 0.005 mm?
Answer #1

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 18.005
std deviation   =σ= 0.0040
probability = P(17.987<X<18) = P(-4.5<Z<-1.25)= 0.1056-0= 0.1056

b)

probability = P(17.987<X<18.013) = P(-4.5<Z<2)= 0.9772-0= 0.9772

c)

therefore corresponding value=mean+z*std deviation= 18.012

d-a)

probability = P(17.987<X<18) = P(-3.6<Z<-1)= 0.1587-0.0002= 0.1585

d-b)

probability = P(17.987<X<18.013) = P(-3.6<Z<1.6)= 0.9452-0.0002= 0.9450

d-c)

corresponding value=mean+z*std deviation= 18.013